Change 4x (x - 2) -1 = 0 into standard form
Found 2 solutions by ikleyn, Clanther:
Answer by ikleyn(52778)

.

4x (x - 2) -1 = 0   in the standard form is


4x^2 - 8x - 1 = 0.            ANSWER



Answer by Clanther(6)

Quadratic equation in standard form ax²+bx+c=0; a is not equal to 0
Use the distributive property to simplify the mathematical expression
4x(x-2)-1=0
4x²-8x-1=0 - standard form